# A Graph Laplacian Eigenvector-based Pre-training Method for Graph Neural Networks

### v2

@@ 1 Introduction
Graph Neural Networks (GNNs) have become a powerful tool in node and graph representation learning, with successful applications across domains ranging from biomedicine (Cantürk et al., 2023; Yan et al., 2024) to social networks (Fan et al., 2019). More recently, graph foundation models (GFMs) are emerging as an exciting field; inspired by the success of large language models (LLMs), researchers are exploring the possibility of creating large graph-based models with emergent capabilities across many domains (Liu et al., 2025; Xie et al., 2022; Wang et al., 2025).
A key ingredient in this effort is the creation of self-supervised tasks which can be performed on large unlabeled graph datasets (Liu et al., 2022). A variety of pre-training methods have been proposed, with the majority of such methods being feature-based, including contrastive losses and feature reconstruction (Liu et al., 2025; Xie et al., 2022). A few structure-based methods, which precompute labels based on graph topology, have been proposed (Peng et al., 2020a; Hwang et al., 2020). However, this category of pre-training approach, recently termed graph property prediction (Liu et al., 2025), remains under-explored. Furthermore, limitations of GNNs in capturing long-range structural information caused by oversmoothing and underreaching (Cai & Wang, 2020; Oono & Suzuki, 2019; Keriven, 2022) pose challenges when approaching graph property prediction tasks which rely on global structure.
@@ 3.2 Graph-level MLP
Eigenvector-learning methods for other applications (such as spectral clustering) use a node-wise MLP head, processing each node’s eigencoordinates independently based on their learned hidden embedding (Shaham et al., 2018; Dwivedi et al., 2021; Cantürk et al., 2023). Such methods fail to address oversmoothing as a node-wise MLP cannot learn long-range interactions; instead, one must use several layers of message passing within the base GNN (Cantürk et al., 2023).
@@ 3.3 Eigenvector prediction
To ensure the model does not output $k$ copies of the trivial eigenvector, we impose orthogonality on the final outputs of the model via QR decomposition, as proposed by Shaham et al. (2018). Let each $\hat{{{\mathbf{u}}}}_{i}$ denote the $i$th predicted eigenvector via LELM after orthogonality has been imposed.
Energy loss, used in Shaham et al. (2018); Dwivedi et al. (2021); Ma & Zhan (2023), computes the sum of Rayleigh quotients:
@@
Energy loss is minimized when the predicted eigenvectors span the same subspace as the ground-truth eigenvectors; to enforce a strict ordering and basis on the predicted eigenvectors, we additionally impose eigenvector loss, used in Mishne et al. (2019):
@@ 4 Experimental Results
Comparison against baseline models: We pre-train a standard Graph Isomorphism Network (GIN) (Xu et al., 2019) and GPS (Rampášek et al., 2022), a graph transformer, using LELM. Once the model has been pre-trained, we replace the graph-level MLP head with a downstream prediction MLP and fine-tune model weights. We evaluate our pre-training framework on three molecular datasets ZINC, ZINC-12k (Sterling & Irwin, 2015) and QM9 (Ramakrishnan et al., 2014). For each of these models, we compare LELM against the same GNN model without pre-training. For both models, pre-training improves performance for all but one of the downstream targets. We record results of our experiments in Table 1.
@@ Appendix A Related works: GNN pre-training methods
Towards the goal of improving graph foundation models, a variety of self-supervised graph pre-training tasks have been proposed. According to the taxonomy provided by Liu et al. (2025); Xie et al. (2022), existing graph pre-training methods can be categorized into two broad categories: contrastive and predictive methods.
Contrastive methods maximize mutual information between pairs of data views using objectives like Jensen-Shannon estimator (Nowozin et al., 2016) or InfoNCE (Oord et al., 2018). Methods can be categorized by the types of views used: graph-node (Sun et al., ; Veličković et al., ; Peng et al., 2020b), subgraph-node (Hu et al., ; Jiao et al., 2020), and subgraph-subgraph Qiu et al. (2020). Some methods also employ graph augmentation to generate two views (You et al., 2020).
Predictive methods, also referred to as generative methods (Liu et al., 2025), self-generate labels and train to predict these labels. A first class of predictive models uses graph reconstruction, whether by using node/edge masking (Xie et al., 2020; Batson & Royer, 2019; Hu et al., ) or using autoencoders (Wang et al., 2017; Kipf & Welling, 2016). A second class of predictive methods are property prediction methods, which precompute underlying graph properties as labels. Examples include statistical properties such as k-hop connectivity (Peng et al., 2020a) or topological properties like a meta-path (Hwang et al., 2020). Overall, there are a lack of works on property prediction-based methods, with the majority of predictive pre-training methods falling under the former category of graph reconstruction (Liu et al., 2025). Our method, LELM, is the first property prediction method to use the graph Laplacian eigenvectors as a pre-training target.
@@ Appendix B Graph Laplacian eigenvector applications
On any bounded-degree $n$-vertex planar graph, the optimal Fiedler cut has ratio $O(\frac{1}{n})$ (Spielman & Teng, 1996).
Positional encodings: The low-frequency Laplacian eigenvectors naturally encode a global position on the graph. As a result, Laplacian positional encodings (LapPE) have been used as a standard positional encoding for graph transformer models (Dwivedi & Bresson, 2020; Rampášek et al., 2022). In practice, directly using the Laplacian eigenvectors as positional encodings creates sign and basis ambiguity issues, as $\psi_{i}$ is an eigenvector of $L\iff-\psi_{i}$ is an eigenvector of $L$. Approaches to solving this problem include designing an architecture which processes the Laplacian eigenvectors in a sign- and basis-invariant manner (Lim et al., 2022) or defining canonical directions for the eigenvectors (Ma et al., 2023).
Spectral GNNs: A variety of methods, known as spectral graph neural networks, use the Laplacian eigendecomposition to learn filters in signal domain (Bo et al., 2023b). Some methods explicitly compute or approximate the $k$ lowest-frequency Laplacian eigenvectors, learning advanced filters on the corresponding eigenvalues (Bruna et al., 2013; Liao et al., 2019; Bo et al., 2023a). Other methods instead learn polynomial filters on the graph (Defferrard et al., 2016; He et al., 2022), circumventing the expensive process of eigendecomposition by learning a $k$-degree polynomial function with respect to $L$, i.e. $p(L)=\theta_{0}I+\theta_{1}L+\dots+\theta_{k}L^{k}$.
Spectral clustering: The Laplacian eigenvectors have also been used for clustering applications. Given a set of data $x_{1},\dots x_{n}$, Belkin & Niyogi (2001) construct a weighted graph $G$ with $n$ nodes using a heat kernel. Then to generate a $k$-dimensional embedding, Belkin & Niyogi (2001) compute the $k$-lowest eigenvectors $\psi_{2},\dots\psi_{k+1}$ (omitting the trivial eigenvector) of the graph Laplacian and assign data point $x_{i}$ the embedding $(\psi_{2}(i),\psi_{3}(i),\dots,\psi_{k+1}(i))$. Shaham et al. (2018); Chen et al. (2022) learn this spectral map using a neural network, allowing for a natural extension of this map to new datapoints.
@@ Appendix D Loss function
Energy loss, used by Shaham et al. (2018); Dwivedi et al. (2021); Ma & Zhan (2023), aims to minimize the sum of Rayleigh quotients:
@@
Eigenvector loss, used by Mishne et al. (2019), measures the difference between each $L\hat{u}_{i}$ and $\lambda_{i}\hat{u}_{i}$:
@@ E.1 Orthogonality
Forced orthogonality, used in Shaham et al. (2018), imposes orthogonality on the final outputs of the model via QR decomposition. In other words, if $\hat{U}^{\prime}$ is the initial output to the model, $Q$ is an $n\times k$ matrix with orthonormal columns, and $R$ is a $k\times k$ upper triangular matrix, then we achieve the final output $\hat{U}$ as such:
@@
Orthogonality loss, used in Dwivedi et al. (2021); Ma & Zhan (2023); Mishne et al. (2019) imposes a softer constraint, encouraging orthogonality by penalizing the model for producing pairwise similar vectors. This can be written as:
@@ Appendix H Alternative structural pre-training targets
  • RWSE: A node-level label computing self-walk probabilities at varying step counts for the diffusion operator (Dwivedi et al., 2021). In our experiments, we use step counts from the interval $[2,22]$.
@@ I.1 Summary
We augment the existing molecular pre-training methods proposed by Hu et al. (2019) with eigenvector-learning. In particular, Hu et al. (2019) propose node-level pre-training tasks (context prediction and masking) on ZINC15 (Sterling & Irwin, 2015), followed by a graph-level supervised pre-training task on ChEMBL (Mayr et al., 2018; Gaulton et al., 2012). We augment the graph-level supervised pre-training step by adding an additional MLP head to predict eigenvectors, and we evaluate on five downstream datasets based on work by Sun et al. (2022).
@@
Table 4: Test ROC-AUC (%, $\uparrow$) performance on 5 molecular prediction tasks when augmenting an existing pre-training method on a GIN base model. Sup. refers to the original supervised pre-training as implemented by Hu et al. (2019), and Sup.+ refers to supervised training with LELM. Results for no pre-training are taken directly from Sun et al. (2022). All methods are tuned over seven learning rates and averaged over three seeds.
@@ I.2 Learning rate tuning
We keep the majority of the settings from Hu et al. (2019) the same. For downstream fine-tuning, we tune over 7 learning rates for fair comparison according to Sun et al. (2022). We run each method and learning rate over 3 seeds, and select the learning rate based on mean validation accuracy over all learning rates.
@@ I.3 Downstream datasets
We briefly list and cite the five downstream datasets here for reference. The five datasets are the datasets chosen in Sun et al. (2022), and are a subset of the eight primary downstream datasets evaluated in Hu et al. (2019).
  • BACE: Qualitative binding results Subramanian et al. (2016)
  • BBBP: Blood-brain barrier penetration Martins et al. (2012)
  • Tox21: Toxicity data Mayr et al. (2016)
  • Toxcast: Toxicology measurements Richard et al. (2016)
  • SIDER: Database of adverse drug reactions (ADR) Kuhn et al. (2016)

### v3

@@ 1 Introduction
Graph Neural Networks (GNNs) have become a powerful tool in node and graph representation learning, with successful applications across domains ranging from biomedicine (canturk2023graph; yan2024empowering) to social networks (fan2019graph). More recently, graph foundation models (GFMs) are emerging as an exciting field; inspired by the success of large language models (LLMs), researchers are exploring the possibility of creating large graph-based models with emergent capabilities across many domains (liu2025graph; xie2022self; wang2025graph).
A key ingredient in this effort is the creation of self-supervised tasks which can be performed on large unlabeled graph datasets (liu2022graph). A variety of pre-training methods have been proposed, with the majority of such methods being feature-based, including contrastive losses and feature reconstruction (liu2025graph; xie2022self). A few structure-based methods, which precompute labels based on graph topology, have been proposed (peng2020self; hwang2020self). However, this category of pre-training approach, recently termed graph property prediction (liu2025graph), remains under-explored. Furthermore, limitations of GNNs in capturing long-range structural information caused by oversmoothing and underreaching (cai2020note; oono2019graph; keriven2022not) pose challenges when approaching graph property prediction tasks which rely on global structure.
@@ 3.2 Graph-level MLP
Eigenvector-learning methods for other applications (such as spectral clustering) use a node-wise MLP head, processing each node’s eigencoordinates independently based on their learned hidden embedding (shaham2018spectralnet; dwivedi2021graph; canturk2023graph). Such methods fail to address oversmoothing as a node-wise MLP cannot learn long-range interactions; instead, one must use several layers of message passing within the base GNN (canturk2023graph).
@@ 3.3 Eigenvector prediction
To ensure the model does not output $k$ copies of the trivial eigenvector, we impose orthogonality on the final outputs of the model via QR decomposition, as proposed by shaham2018spectralnet. Let each $\hat{{{\mathbf{u}}}}_{i}$ denote the $i$th predicted eigenvector via LELM after orthogonality has been imposed.
Energy loss, used in shaham2018spectralnet; dwivedi2021graph; ma2023self, computes the sum of Rayleigh quotients:
@@
Energy loss is minimized when the predicted eigenvectors span the same subspace as the ground-truth eigenvectors; to enforce a strict ordering and basis on the predicted eigenvectors, we additionally impose eigenvector loss, used in mishne2019diffusion:
@@ 4 Experimental Results
Comparison against baseline models: We pre-train a standard Graph Isomorphism Network (GIN) (GIN) and GPS (rampavsek2022recipe), a graph transformer, using LELM. Once the model has been pre-trained, we replace the graph-level MLP head with a downstream prediction MLP and fine-tune model weights. We evaluate our pre-training framework on three molecular datasets ZINC, ZINC-12k (sterling2015zinc) and QM9 (ramakrishnan2014quantum). For each of these models, we compare LELM against the same GNN model without pre-training. For both models, pre-training improves performance for all but one of the downstream targets. We record results of our experiments in Table 1.
@@ Appendix A Related works: GNN pre-training methods
Towards the goal of improving graph foundation models, a variety of self-supervised graph pre-training tasks have been proposed. According to the taxonomy provided by liu2025graph; xie2022self, existing graph pre-training methods can be categorized into two broad categories: contrastive and predictive methods.
Contrastive methods maximize mutual information between pairs of data views using objectives like Jensen-Shannon estimator (nowozin2016f) or InfoNCE (oord2018representation). Methods can be categorized by the types of views used: graph-node (suninfograph; velivckovicdeep; peng2020graph), subgraph-node (hustrategies; jiao2020sub), and subgraph-subgraph qiu2020gcc. Some methods also employ graph augmentation to generate two views (you2020graph).
Predictive methods, also referred to as generative methods (liu2025graph), self-generate labels and train to predict these labels. A first class of predictive models uses graph reconstruction, whether by using node/edge masking (xie2020noise2same; batson2019noise2self; hustrategies) or using autoencoders (wang2017mgae; kipf2016variational). A second class of predictive methods are property prediction methods, which precompute underlying graph properties as labels. Examples include statistical properties such as k-hop connectivity (peng2020self) or topological properties like a meta-path (hwang2020self). Overall, there are a lack of works on property prediction-based methods, with the majority of predictive pre-training methods falling under the former category of graph reconstruction (liu2025graph). Our method, LELM, is the first property prediction method to use the graph Laplacian eigenvectors as a pre-training target.
@@ Appendix B Graph Laplacian eigenvector applications
On any bounded-degree $n$-vertex planar graph, the optimal Fiedler cut has ratio $O(\frac{1}{n})$ (spielman1996spectral).
Positional encodings: The low-frequency Laplacian eigenvectors naturally encode a global position on the graph. As a result, Laplacian positional encodings (LapPE) have been used as a standard positional encoding for graph transformer models (dwivedi2020generalization; rampavsek2022recipe). In practice, directly using the Laplacian eigenvectors as positional encodings creates sign and basis ambiguity issues, as $\psi_{i}$ is an eigenvector of $L\iff-\psi_{i}$ is an eigenvector of $L$. Approaches to solving this problem include designing an architecture which processes the Laplacian eigenvectors in a sign- and basis-invariant manner (lim2022sign) or defining canonical directions for the eigenvectors (ma2023laplacian).
Spectral GNNs: A variety of methods, known as spectral graph neural networks, use the Laplacian eigendecomposition to learn filters in signal domain (bo2023survey). Some methods explicitly compute or approximate the $k$ lowest-frequency Laplacian eigenvectors, learning advanced filters on the corresponding eigenvalues (bruna2013spectral; liao2019lanczosnet; bo2023specformer). Other methods instead learn polynomial filters on the graph (defferrard2016convolutional; he2022convolutional), circumventing the expensive process of eigendecomposition by learning a $k$-degree polynomial function with respect to $L$, i.e. $p(L)=\theta_{0}I+\theta_{1}L+\dots+\theta_{k}L^{k}$.
Spectral clustering: The Laplacian eigenvectors have also been used for clustering applications. Given a set of data $x_{1},\dots x_{n}$, belkin2001laplacian construct a weighted graph $G$ with $n$ nodes using a heat kernel. Then to generate a $k$-dimensional embedding, belkin2001laplacian compute the $k$-lowest eigenvectors $\psi_{2},\dots\psi_{k+1}$ (omitting the trivial eigenvector) of the graph Laplacian and assign data point $x_{i}$ the embedding $(\psi_{2}(i),\psi_{3}(i),\dots,\psi_{k+1}(i))$. shaham2018spectralnet; chen2022specnet2 learn this spectral map using a neural network, allowing for a natural extension of this map to new datapoints.
@@ Appendix D Loss function
Energy loss, used by shaham2018spectralnet; dwivedi2021graph; ma2023self, aims to minimize the sum of Rayleigh quotients:
@@
Eigenvector loss, used by mishne2019diffusion, measures the difference between each $L\hat{u}_{i}$ and $\lambda_{i}\hat{u}_{i}$:
@@ E.1 Orthogonality
Forced orthogonality, used in shaham2018spectralnet, imposes orthogonality on the final outputs of the model via QR decomposition. In other words, if $\hat{U}^{\prime}$ is the initial output to the model, $Q$ is an $n\times k$ matrix with orthonormal columns, and $R$ is a $k\times k$ upper triangular matrix, then we achieve the final output $\hat{U}$ as such:
@@
Orthogonality loss, used in dwivedi2021graph; ma2023self; mishne2019diffusion imposes a softer constraint, encouraging orthogonality by penalizing the model for producing pairwise similar vectors. This can be written as:
@@ Appendix H Alternative structural pre-training targets
  • RWSE: A node-level label computing self-walk probabilities at varying step counts for the diffusion operator (dwivedi2021graph). In our experiments, we use step counts from the interval $[2,22]$.
@@ I.1 Summary
We augment the existing molecular pre-training methods proposed by hu2019strategies with eigenvector-learning. In particular, hu2019strategies propose node-level pre-training tasks (context prediction and masking) on ZINC15 (sterling2015zinc), followed by a graph-level supervised pre-training task on ChEMBL (mayr2018large; gaulton2012chembl). We augment the graph-level supervised pre-training step by adding an additional MLP head to predict eigenvectors, and we evaluate on five downstream datasets based on work by sun2022does.
@@
Table 4: Test ROC-AUC (%, $\uparrow$) performance on 5 molecular prediction tasks when augmenting an existing pre-training method on a GIN base model. Sup. refers to the original supervised pre-training as implemented by hu2019strategies, and Sup.+ refers to supervised training with LELM. Results for no pre-training are taken directly from sun2022does. All methods are tuned over seven learning rates and averaged over three seeds.
@@ I.2 Learning rate tuning
We keep the majority of the settings from hu2019strategies the same. For downstream fine-tuning, we tune over 7 learning rates for fair comparison according to sun2022does. We run each method and learning rate over 3 seeds, and select the learning rate based on mean validation accuracy over all learning rates.
@@ I.3 Downstream datasets
We briefly list and cite the five downstream datasets here for reference. The five datasets are the datasets chosen in sun2022does, and are a subset of the eight primary downstream datasets evaluated in hu2019strategies.
  • BACE: Qualitative binding results subramanian2016computational
  • BBBP: Blood-brain barrier penetration martins2012bayesian
  • Tox21: Toxicity data mayr2016deeptox
  • Toxcast: Toxicology measurements richard2016toxcast
  • SIDER: Database of adverse drug reactions (ADR) kuhn2016sider